OpenStudy (anonymous):

A ball is dropped from a treetop 256 feet above the ground. How long does it take to hit the ground? (s = 16t^2)

OpenStudy (dumbcow):

solve equation for t t = sqrt(s/16) = sqrt(s)/4 t = sqrt(256)/4 t = 16/4 t=4